# User-customizable Shared Control for Fine Teleoperation via Virtual Reality

### v1

@@ I Introduction
While shared control alleviates excess workload exerted on operators, it may also give rise to new issues that impact task performance and impede dexterity training [5]. One prominent example of such an issue is model misalignment. This phenomenon occurs when the division in control between the human and autonomy creates a misunderstanding in how the operator expects the robot to behave [6]. In direct teleoperation, model misalignment may arise whenever a robot has higher degrees-of-freedom (DoF) than the user’s control interface [7]. Despite this interface asymmetry in DoF, a user with adequate training will still gradually build a mental model of the robot’s behavior. In contrast, the extra layer of arbitration introduced in shared control exacerbates the user’s difficulty in comprehending the system. A communication medium for user feedback is therefore essential to how an operator understands the arbitration [4].
@@
Aside from equipping operators with suitable feedback, another integral aspect of shared control is user customizability. Over a long-term interaction, operators will experience variations in their capabilities, preferences, and environments [8]. Shared control must account for these changes and adapt accordingly, especially in terms of how control authority is assigned. Prior works have developed promising frameworks for adaptive shared control by enabling users to either interactively customize the parameters that characterize the arbitration function [9] or suggest corrections for the resulting robot behavior [10, 11]. However, these works rely entirely on the robot’s legible state to convey the effects of a user’s modifications to the arbitration procedure.
@@
Figure 1: Illustration of the Kinova Gen 3 robotic arm being teleoperated to play the buzz wire game by a user wearing a virtual reality (VR) headset. The VR controller held by the operator will be represented as a handle in the virtual world, where its simulated effects also have a direct consequence on the physical counterpart played by the robot.
@@
In this paper, we establish user-customizable shared control by directly communicating the internal parameters of the arbitration function to an operator for refinement. By unveiling these parameters, users can better understand the arbitration process and thus effectively edit its inner workings for personalized outcomes. We ground this idea in a teleoperation task, where operators must remotely control a 7-DoF robotic arm in virtual reality (VR) to complete the popular buzz wire game (see Fig. 1). The buzz wire game is a suitable testbed due to the need for precise control over both translational and rotational robot motion. This game also requires significant eye-hand coordination and sustained operator focus [12, 13]. Furthermore, the game is frequently utilized for skill training in domains that would benefit from robot teleoperation, such as surgery [14]. To assess the quality and longevity of the proposed shared control method, we conducted a user study that explores patterns in participant performance over repeated interactions.
@@ II Related Work
Sensory feedback can be supplied across multiple modalities in teleoperation. For example, haptic devices are widely used in bilateral teleoperation to simultaneously enhance an operator’s task efficacy and encourage growth in skill [4, 2]. Other modalities, e.g., audio, force, and visual feedback, have enjoyed similar success [15, 16, 17]. In recent years, VR headsets have proven advantageous in complex manipulation tasks where high-DoF robots are remotely controlled [18, 19, 20] . Moreover, VR interfaces offer operators the opportunity to modify the inner workings of the teleoperated robot, which is an asset we exploit in this work.
In addition to communication, shared control must also consider arbitration, i.e., how to “blend” a user’s inputs with the autonomous controller’s outputs [4]. An arbitration function is typically described by a set of parameters that determine how control is allocated. These parameters could be fixed, e.g., to primarily favor the human leading, or they could be dynamically updated. In the latter case, a vast bulk of the shared control literature resorts to heuristics for parameter updates, such as safety, fluency, agreement, and confidence [21, 22, 23]. User-personalized parameters for shared control can also be configured by balancing out the operator’s capability to independently complete a teleoperation task with their need for autonomous assistance [24, 25]. However, the user is unable to configure arbitration parameters in these approaches if no bidirectional communication is available.
Heuristics-based arbitration may be suitable for certain users, but it could also lead to reduced satisfaction or performance in others. To identify the exact form of the “optimal” arbitration function and its corresponding parameters is an intractable problem in real-world teleoperation [9]. Therefore, it may be more appropriate to empower users with the ability to customize the arbitration parameters online. While previous approaches to customizable shared control have utilized verbal feedback [9, 26] or corrections [10, 11] to grant this ability, they solely depended on the robot’s observable behavior to explain the control-sharing. Relying on the robot’s appearance and movement to communicate any underlying autonomous assistance is often insufficient [27]. Instead, our paper opts for greater transparency by externalizing the factors that influence arbitration and making them editable to operators.
@@ III User-customizable Shared Control
III-A Formulation
The shared control system considered in this work is driven by the following interaction data. At time $t$, a human operator’s measured state, $\mathbf{x}_{h}(t)\in\mathbb{R}^{n_{h}}$ (e.g., joystick deflections), is fed into a control interface to produce control commands, $\mathbf{u}_{h}(t)\in\mathbb{R}^{m_{h}}$, for direct robot teleoperation. Similarly, a robot’s state and the autonomy’s assistive commands can be denoted as $\mathbf{x}_{r}(t)\in\mathbb{R}^{n_{r}}$ (e.g., joint angles, Cartesian poses) and $\mathbf{u}_{r}(t)\in\mathbb{R}^{m_{r}}$, respectively.
Under this setup, the output commands of a shared control system adhere to an arbitration process:
$\mathbf{u}_{sc}(t)=\bm{\beta}_{\bm{\theta}}(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t)),$ (1)
where $\bm{\beta}(\cdot)$ is an arbitration function parameterized by time-varying, $\bm{\theta}(t)$, or time-invariant variables, $\bm{\theta}$. As an example, the standard linear blending scheme could be written as:
@@ III-A Formulation
where $\alpha_{\bm{\theta}}\in[0,1]$ is the blending variable and is also a function parameterized by $\bm{\theta}$ [9]. Note that $\alpha$ is frequently treated as a scalar, but in our approach we will adopt a matrix representation to facilitate more granular authority over the system behavior. Let $\mathbf{A}_{\bm{\theta}}=\text{diag}(\alpha_{1},\ldots,\alpha_{m_{r}})$ be a positive definite diagonal arbitration matrix, with $\forall i,\alpha_{i}\in[0,1]$, such that the blending scheme is now:
@@
where $\mathbf{I}$ is an identity matrix.
Selecting the parameters, $\bm{\theta}$, of an arbitration function, $\bm{\beta}_{\bm{\theta}}(\cdot)$, is critical in shaping the robot’s assistive behavior. Instead of manually selecting a fixed choice of $\bm{\theta}$, an “optimal” selection is often derived by solving an optimization problem with an objective function, $\Gamma$, such as:
$\Gamma=\mathcal{L}_{T}(\mathbf{x}_{r}(t))+\sum_{t=0}^{T-1}\mathcal{L}\left(\mathbf{x}_{r}(t),\bm{\beta}_{\bm{\theta}}\left(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t)\right)\right),$ (4)
where $\mathcal{L}_{T}(\cdot)$ is a terminal cost for a horizon, $T$ (e.g., error between the intended and actual target robot states), and $\mathcal{L}(\cdot)$ is an incremental cost (e.g., human effort). Finding a time-invariant $\bm{\theta}$ to minimize $\Gamma$ will ignore the dynamic nature of the human-robot collaboration. Even if a solution is found with time-varying $\bm{\theta}(t)$, the core issue remains that $\Gamma$ might not be the true cost function of the shared control, $\Gamma^{*}$, nor would a tractable solution to $\Gamma^{*}$ necessarily exist [9].
@@ III-B Feedback-informed User Optimization
Figure 2: Diagram of the user-customizable shared control architecture.
To circumvent the intractable problem of identifying $\Gamma^{*}$ and its “optimal” parameters, we present a method that allows users to personalize the arbitration function by configuring their preferred $\bm{\theta}$. The key insight behind our formulation is to establish a bidirectional communication channel between the human and robot. We formalize this channel as a vector-valued function $\bm{\psi}(\cdot)$ (dropping $t$ for brevity):
@@
where the robot administers feedback, $\mathbf{v}_{r}$, to the operator about the shared control, while the user performs parameter updates, $\Delta\bm{\theta}$, via an interface action, $v_{h}$.
For instance, a scalar blending quantity, $\alpha$, could be derived from the following piece-wise function [9, 25]:
$\alpha$ =$=\begin{cases}0&C\leq\theta_{1},\\ \frac{\theta_{3}}{\theta_{2}-\theta_{1}}\cdot C&\theta_{1}<C\leq\theta_{2},\\ \theta_{3}&C>\theta_{2},\end{cases}$ (6)
$\text{where}\quad\theta_{j}$ $=\psi_{j}(\mathbf{v}_{r},v_{h}),\quad j=1:K.$
The $\delta\theta_{j}$ in the $K$ parameter mappings, $\psi_{j}(\mathbf{v}_{r},v_{h})=\theta_{j}\pm\delta\theta_{j}$, could be obtained from discrete verbal commands made by the operator, $v_{h}\in\mathbb{Z}^{+}$ (e.g., “more/less assistance”). Here, the autonomy’s confidence in its estimation of a user’s goal, $C$, is independent of the resulting arbitration level. We highlight that most previous efforts at user customization of $\theta_{j}$ have solely depended on the robot’s legible state, $\mathbf{v}_{r}=\mathbf{x}_{r}$, for operator feedback [9, 26].
In the multi-dimensional arbitration scenario, one could choose to represent $\mathbf{A}_{\bm{\theta}}$ in a more general form:
@@
$\text{where}\quad\alpha_{i}$ $=\sum_{j=1}^{K}w_{j}\psi_{j}(\mathbf{v}_{r},v_{h}),\quad i=1,\ldots,m_{r}.$
Each diagonal element, $\alpha_{i}$, is governed by potentially more than one factor, $\theta_{j}$, resultant from $\psi_{j}(\mathbf{v}_{r},v_{h})$, as determined by the selection weight, $w_{j}$. Whether the arbitration is expressed as Eq. 6 or Eq. 7, the aspiration behind permitting operator customization of $\bm{\theta}$ is to converge on a user-elected optimum, without concretely specifying $\Gamma^{*}$. The overall framework architecture is shown in Fig. 2.
As the feedback variables, $\mathbf{v}_{r}$ and $v_{h}$, dramatically impact how users customize the shared control behavior, we suggest two improvements over prior works in how to produce these variables. First, the robot’s communication flow should illuminate more than its external state, $\mathbf{x}_{r}$. In Section IV, we demonstrate how a set of human-interpretable signals, $\mathcal{V}$ (i.e., visualizations in VR), can be incorporated into $\mathbf{v}_{r}=\{\mathbf{x}_{r},\mathcal{V}\}$ to help operators tune $\bm{\theta}$. Second, an operator’s parameter updates, $\Delta\bm{\theta}$, should have a proportionate effect on robot behavior. This proportionality can be achieved by letting the operator submit actions, $v_{h}\in[0,1]$ (e.g., using a graphical slider in VR), over the same constrained range as the arbitration parameters, $\forall j,\theta_{j}\in[0,1]$.
@@ IV-A Testbed: Buzz Wire Game
In the buzz wire game, the operator’s goal is to carefully guide a loop handle in $\mathrm{SE}(3)$ from start to end while minimizing collisions with the wire. The game is a well-known assessment environment for fine motor skill and eye-hand coordination [12, 13]. Moreover, 3D buzz wire games are advantageous when evaluating dexterity tasks that benefit from stereoscopic depth perception [12]. As precise motor control, eye-hand coordination, and 3D spatial awareness are critical capabilities for teleoperation in general, we deem this game to be a pertinent testbed. Immersive VR versions of the game have also been validated as an effective simulation for operator training, e.g., in stroke rehabilitation [28, 29]. Hence, we believe a VR setup of the 3D game serves as a suitable experimental arena for our work.
Fig. 1 demonstrates our buzz wire game replica for robot teleoperation. The layout is comprised of a Kinova Gen 3 arm with a loop handle extension at its end-effector to complete the actual game, as well as a person wearing an HTC Vive headset to perform the same task in VR with a handheld controller. A model-based perspective of the buzz wire game and the robot are also constructed in the VR scene. An electric circuit board imitated the game’s “buzz” feature by signaling collisions to the operator in VR. Multimodal feedback cues were triggered upon collision, with red lights flashing and vibrations in the handheld input device.
@@ IV-B Shared Assistance in Teleoperation
where $\mathbf{M}\in\mathbb{R}^{6\times 6}$, $\mathbf{k}\in\mathbb{R}^{6}$, and $\mathbf{d}\in\mathbb{R}^{6}$ are the pre-defined mass, stiffness, and damping terms respectively, while $\mathbf{f}_{\mathrm{ext}}\in\mathbb{R}^{6}$ is the externally applied force, $\mathbf{e}\in\mathbb{R}^{6}$ is the pose error vector, and $\odot$ denotes component-wise multiplication. In our buzz wire setup, the pose error is the difference between the handheld VR controller’s pose, $\mathbf{x}_{h}\in\mathbb{R}^{6}$, and the robot’s state, ${\mathbf{x}_{r}\in\mathbb{R}^{6}}$.
The robot autonomy can then be integrated into Eq. 8 as an assistive force. By reformulating Eq. 8 as the linear blending scheme shown in Eq. 3, we establish an admittance control policy for shared control as follows:
@@
In this formulation, $\mathbf{f}_{\mathrm{ext}}$ has been replaced by a wrench vector, $\mathbf{w}_{a}\in\mathbb{R}^{6}$, to represent the autonomous assistance in both translational and rotational space.
@@ IV-C Real-time Assistive Wrench
Figure 3: Illustration of how the attractive and repulsive forces are computed given a control point $\mathbf{p}_{i}$ (dark blue) and environment point $\mathbf{p}_{k}$ (red).
Generating real-time autonomous assistance is vital for teleoperation to ensure a reactive user experience and smooth robot motion. To fulfil this requirement, we efficiently compute the assistive wrench, $\mathbf{w}_{a}$, using potential fields, where targets produce attractive vectors and obstacles yield repulsive vectors. Attractive and repulsive vectors are typically calculated between a pair of control and environment points [30]. In our setting, we define eight control points as the vertices of an octagon that encloses the circular end-effector. The environment points are acquired in real-time given point cloud data detected using two depth cameras.
where each function-shaping parameter ${\lambda_{i,k}\in\{\lambda_{\mathrm{rep}},\lambda_{\mathrm{att}}\}}$, ${d_{i,k}\in\{d_{\mathrm{rep}},d_{\mathrm{att}}\}}$, ${\hat{\mathbf{d}}_{i,k}=\pm\frac{\mathbf{d}_{i,k}}{\|\mathbf{d}_{i,k}\|}}$ has two possible values depending on the relationship between control point $i$ and environment point $k$ (repulsive or attractive). Hence, forces $\mathbf{f}_{i,k}$ can either be repulsive or attractive, with $f_{\mathrm{max}}$ determining the maximum magnitude of the generated force and $\rho$ adjusting the change rate in force given distance.
@@
$\bm{\tau}_{i,k}=\mathbf{f}_{i,k}\times(\mathbf{p}_{\mathrm{com}}-\mathbf{p}_{i})$ (13)
where $\mathbf{p}_{\mathrm{com}}$ is the center of mass of the robot’s end-effector. Fig. 3 illustrates an example of how attractive forces and repulsive torques are calculated for the buzz wire game given a pair of control and environment points $(\mathbf{p}_{i},\mathbf{p}_{k})$.
The resultant assistance wrench, $\mathbf{w}_{a}=\begin{bmatrix}\mathbf{f}_{\mathrm{net}}^{\mathsf{T}}&\bm{\tau}_{\mathrm{net}}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}$, is obtained by aggregating a force component, ${\mathbf{f}_{\mathrm{net}}\in\mathbb{R}^{3}}$, and a torque component, $\bm{\tau}_{\mathrm{net}}\in\mathbb{R}^{3}$, each calculated as:
@@
where ${\mathbf{f}_{i}=\sum_{k}\mathbf{f}_{i,k}}$, ${\bm{\tau}_{i}=\sum_{k}\bm{\tau}_{i,k}}$, $i\in\{1,\ldots,8\}$, and $k\in\{1,\ldots,N\}$.

### v2

# User-customizable Shared Control for Robot Teleoperation via Virtual Reality

@@ I Introduction
While shared control alleviates excess workload exerted on operators, it may also give rise to new issues that impact task performance and impede dexterity training [5]. One prominent example of such an issue is model misalignment. This phenomenon occurs when the division in control between the human and autonomy creates a misunderstanding in how the operator expects the robot to behave [6]. In direct teleoperation, model misalignment may arise whenever a robot has higher degrees-of-freedom (DoF) than the user’s control interface [7]. Despite this interface asymmetry in DoF, a user with adequate training should still gradually build a mental model of the robot’s behavior. In contrast, the extra layer of arbitration introduced in shared control exacerbates the user’s difficulty in comprehending the system. A communication medium for user feedback is therefore essential to how an operator understands the arbitration [4].
@@
Aside from equipping operators with suitable feedback, another integral aspect of shared control is user customizability. Over a long-term interaction, operators will experience variations in their capabilities, preferences, and environments [8]. Shared control must account for these changes and adapt accordingly, especially in terms of how control authority is assigned [9]. Prior works have developed promising frameworks for adaptive shared control by enabling users to interactively customize the parameters that characterize the arbitration function [10, 11] or suggest corrections for the resulting robot behavior [12, 13]. However, these works rely entirely on the robot’s legible state to convey the effects of a user’s modifications to the arbitration procedure.
@@
Figure 1: A user wearing a virtual reality (VR) headset teleoperates the Kinova Gen 3 robotic arm to play the buzz wire game. The VR controller acts as a handle in the virtual world, where its simulated effects also have a direct consequence on the physical counterpart played by the robot.
@@
In this paper, we establish user-customizable shared control by directly communicating the internal arbitration parameters to an operator for refinement. This transparency allows users to better understand the arbitration process and thus effectively modify it for personalized outcomes. We ground this idea in a teleoperation task where operators must control a 7-DoF robotic arm in virtual reality (VR) to complete the buzz wire game (see Fig. 1). The buzz wire game is a suitable testbed for fine robot control, as it requires both translational and rotational motion, significant eye-hand coordination, and sustained operator focus [14, 15]. Moreover, the game is frequently utilized for skill training in domains that would benefit from robot teleoperation, such as surgery [16]. To assess the quality and longevity of the proposed shared control method, we conducted a user study to analyze how participant performance evolves over repeated interactions.
@@ II Related Work
Sensory feedback can be supplied across multiple modalities in teleoperation. For example, haptic devices are widely used in bilateral teleoperation to simultaneously enhance an operator’s task efficiency and encourage growth in skill [4, 2]. Other modalities, e.g., audio, force, and visual feedback, have enjoyed similar success [17, 18, 19]. In recent years, VR headsets have proven advantageous in complex manipulation tasks where high-DoF robots are remotely controlled [20, 21, 22]. Furthermore, VR interfaces offer operators the opportunity to modify the inner workings of the teleoperated robot, which is an asset we exploit in this work.
In shared control teleoperation, the challenge is to also effectively “blend” user inputs with an autonomous controller’s outputs. This blending or arbitration process is typically described by parameters that determine how control is allocated. Arbitration parameters may be fixed, e.g., to favor the human leading, or dynamically updated based on heuristics [4], such as safety, fluency, and confidence [23, 24, 25]. These parameters can also be personalized to the user by balancing out their capability to independently complete a task with their need for assistance [26, 27]. However, striking the desired balance in autonomy is an open problem [9], as heuristically tuning arbitration parameters may not be suitable for all users – even when the tuning is personalized.
Therefore, it may be more appropriate to empower users with the ability to customize the arbitration process. While the literature on customizable shared control is sparse, some previous works have successfully granted this ability via user feedback [10, 11] or “corrections” [12, 13]. Hence, users can haptically or verbally correct arbitration parameters on a trial-by-trial basis, e.g., requesting “less assistance” for the next interaction. Although effective, these works relied solely on the robot’s appearance and movement to communicate the underlying autonomous assistance, which is often insufficient at explaining arbitration to users [27, 28]. Instead, our paper opts for greater transparency by externalizing the factors that influence arbitration and making them editable to operators.
@@ III User-customizable Shared Control
III-A Problem Setup
The shared control system considered in this work is driven by the following interaction data. At time $t$, a human operator’s measured state, $\mathbf{x}_{h}(t)\in\mathbb{R}^{n_{h}}$ (e.g., joystick deflections), is fed into a control interface to produce control commands, $\mathbf{u}_{h}(t)\in\mathbb{R}^{m_{h}}$, for direct robot teleoperation. Similarly, a robot’s state and the autonomy’s assistive commands are $\mathbf{x}_{r}(t)\in\mathbb{R}^{n_{r}}$ (e.g., joint angles) and $\mathbf{u}_{r}(t)\in\mathbb{R}^{m_{r}}$, respectively. The output commands of the shared control system, $\mathbf{u}_{sc}(t)=\bm{\beta}_{\bm{\theta}}(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t))$, adhere to an arbitration function, $\bm{\beta}(\cdot,\cdot)$, parameterized by $\bm{\theta}\in\mathbb{R}^{L}$. For instance, a linear blending scheme can be written as:
@@ III-A Problem Setup
where $\alpha_{\bm{\theta}}\in[0,1]$ is the blending variable [10]. While $\alpha$ is usually a scalar, we will adopt a matrix representation for more granular authority over the system behavior. Let $\mathbf{A}_{\bm{\theta}}=\text{diag}(\alpha_{1},\ldots,\alpha_{m_{r}})$ be a positive definite diagonal arbitration matrix, with $\forall i,\alpha_{i}\in[0,1]$, such that linear blending is:
@@
where $\mathbf{I}$ is an identity matrix of appropriate dimensions.
Selecting the parameters, $\bm{\theta}$, of an arbitration function, $\bm{\beta}_{\bm{\theta}}(\cdot,\cdot)$, is critical in shaping the robot’s assistive behavior. Instead of manually selecting a fixed $\bm{\theta}$, an “optimal” selection is often derived by solving an optimization problem with a sensible objective, $\Gamma$, like minimizing human effort. The issue with this approach is that $\Gamma$ may not be the true cost function of the shared control, $\Gamma^{*}$, nor would a tractable solution necessarily exist [10].
@@ III-B Feedback-informed User Optimization
To circumvent the intractable problem of identifying $\Gamma^{*}$ and its “optimal” parameters, we present a method that allows users to customize the arbitration function by configuring their preferred $\bm{\theta}$. The core concept of our method is to establish a bidirectional communication channel between the human and robot. We formalize this channel as a vector-valued function $\bm{\psi}(\cdot,\cdot)$ (dropping $t$ for brevity):
@@
where the robot administers feedback, $\mathbf{v}_{r}$, to the operator about the shared control, while the user performs parameter updates via an interface action, $v_{h}$. For example, parameter updates could be obtained from discrete verbal commands made by the operator, $v_{h}\in\mathbb{Z}^{+}$, e.g., “more/less assistance”. Most previous efforts at user customization of $\bm{\theta}$ have solely depended on the robot’s legible state, $\mathbf{v}_{r}=\mathbf{x}_{r}$, for operator feedback [10, 11].
Figure 2: The user-customizable shared control architecture.
Using this channel formulation to obtain parameters, $\bm{\theta}$, we represent the arbitration matrix, $\mathbf{A}_{\bm{\theta}}$, as follows:
@@
$\text{where}\quad\alpha_{j}$ $=\sum_{l=1}^{L}w_{l}\psi_{l}(\mathbf{v}_{r},v_{h}),\quad j=1,\ldots,m_{r}$
Each scalar blending quantity, $\alpha_{j}$, is governed by potentially more than one parameter, $\theta_{l}$, output as an element of the channel, $\psi_{l}(\mathbf{v}_{r},v_{h})$, depending on the selection weight, $w_{l}$. Our overall framework architecture is shown in Fig. 2.
Since the feedback variables, $\mathbf{v}_{r}$ and $v_{h}$, dramatically impact how users perceive the shared control, we recommend two requirements for generating these variables. First, the robot should communicate more than its external state, $\mathbf{x}_{r}$, to help operators tune $\bm{\theta}$. In Section IV, we demonstrate how a set of human-interpretable cues, $\mathcal{V}$ (i.e., visualizations in VR), can be incorporated into $\mathbf{v}_{r}=\{\mathbf{x}_{r},\mathcal{V}\}$. Second, an operator’s parameter updates should have a proportionate effect on robot behavior. This proportionality can be achieved by letting the operator submit actions, $v_{h}\in[0,1]$ (e.g., using a graphical slider in VR), over the same constrained range as the arbitration parameters, $\forall l,\theta_{l}\in[0,1]$.
@@ IV-A Testbed: Buzz Wire Game
In the buzz wire game, the operator’s goal is to carefully guide a loop handle in $\mathrm{SE}(3)$ from start to end while minimizing collisions with the wire. The game is a well-known assessment environment for fine motor skill and eye-hand coordination [14, 15]. Moreover, 3D buzz wire games are advantageous when evaluating dexterity tasks that benefit from stereoscopic depth perception [14]. As precise motor control, eye-hand coordination, and 3D spatial awareness are critical capabilities for teleoperation in general, we deem this game to be a pertinent testbed. Immersive VR versions of the game have also been validated as an effective simulation for operator training, e.g., in stroke rehabilitation [29].
Fig. 1 demonstrates our buzz wire game replica for robot teleoperation. The layout is comprised of a Kinova Gen 3 arm with a loop handle extension at its end-effector to complete the physical game, as well as a person wearing an HTC Vive headset to perform the same task in VR with a handheld controller. The virtual scene is an environment model reconstruction of the buzz wire game and robot, which has been shown to be less cognitively demanding and more usable in VR contexts than point cloud or image renderings [30]. An electric circuit board on the physical setup also imitates the game’s “buzz” feature by detecting collisions. Collisions are then relayed in VR via flashing red lights and vibrations in the handheld input device.
@@ IV-B Shared Assistance in Teleoperation
where $\mathbf{M}\in\mathbb{R}^{6\times 6}$, $\mathbf{k}\in\mathbb{R}^{6}$, and $\mathbf{d}\in\mathbb{R}^{6}$ are mass, stiffness, and damping terms, respectively, while $\mathbf{f}_{\mathrm{ext}}\in\mathbb{R}^{6}$ is the externally applied force, $\mathbf{e}\in\mathbb{R}^{6}$ is the pose error vector, and $\odot$ denotes component-wise multiplication. In our buzz wire setup, $\mathbf{e}$ is the difference between the handheld VR controller’s pose, $\mathbf{x}_{h}\in\mathbb{R}^{6}$, and the robot’s state, ${\mathbf{x}_{r}\in\mathbb{R}^{6}}$.
Robot autonomy can then be integrated into Eq. 5 as an assistive force by reformulating the linear blending in Eq. 2 to create an admittance control policy for shared control:
@@
Here, $\mathbf{f}_{\mathrm{ext}}$ is replaced by a wrench, $\mathbf{w}_{a}\in\mathbb{R}^{6}$, that represents autonomous assistance in translational and rotational space.
@@ IV-C Real-time Assistive Wrench
Generating real-time autonomous assistance for teleoperation is vital in ensuring reactive and smooth robot motion. To fulfil this requirement, we efficiently compute the assistive wrench, $\mathbf{w}_{a}$, using potential fields, where targets produce attractive vectors and obstacles yield repulsive vectors. Attractive and repulsive vectors are typically calculated between a pair of control and environment points [31]. In our setting, we define eight control points as the vertices of an octagon that encloses the circular end-effector. The environment points are acquired in real-time given point cloud data detected using two depth cameras.
where each function-shaping parameter ${\lambda_{i,k}\in\{\lambda_{\mathrm{rep}},\lambda_{\mathrm{att}}\}}$, ${d_{i,k}\in\{d_{\mathrm{rep}},d_{\mathrm{att}}\}}$, and ${\hat{\mathbf{d}}_{i,k}=\pm\frac{\mathbf{d}_{i,k}}{\|\mathbf{d}_{i,k}\|}}$ has two possible values depending on whether the relationship between $\mathbf{p}_{k}$ and $\mathbf{p}_{i}$ is repulsive or attractive. Hence, forces $\mathbf{f}_{i,k}$ are either repulsive or attractive, with $f_{\mathrm{max}}$ determining the maximum magnitude of the generated force and $\rho$ adjusting the change rate in force given distance.
@@
Figure 3: Illustration of how the attractive and repulsive forces are computed given a control point $\mathbf{p}_{i}$ (dark blue) and environment point $\mathbf{p}_{k}$ (green/red).
@@
$\bm{\tau}_{i,k}=\mathbf{f}_{i,k}\times(\mathbf{p}_{\mathrm{com}}-\mathbf{p}_{i}),$ (10)
where $\mathbf{p}_{\mathrm{com}}$ is the center of mass of the robot’s end-effector. Fig. 3 illustrates an example of how attractive forces and repulsive torques are calculated for the buzz wire game from a pair of control and environment points $(\mathbf{p}_{i},\mathbf{p}_{k})$.
The resultant assistive wrench, $\mathbf{w}_{a}=\begin{bmatrix}\mathbf{f}_{\mathrm{net}}^{\mathsf{T}}&\bm{\tau}_{\mathrm{net}}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}$, is derived by aggregating a force component, ${\mathbf{f}_{\mathrm{net}}\in\mathbb{R}^{3}}$, and a torque component, $\bm{\tau}_{\mathrm{net}}\in\mathbb{R}^{3}$, each calculated as:
@@
where ${\mathbf{f}_{i}=\sum_{k}\mathbf{f}_{i,k}}$ and ${\bm{\tau}_{i}=\sum_{k}\bm{\tau}_{i,k}}$.